OpenStudy (anonymous):

will give a great medal *** An aquarium with a square base has no top. There is a metal frame. Glass costs 3 dollars/m^2 and the frame costs 10 dollars/m. The volume is to be 20 m^3. Express the total cost C in terms of the height h in meters. (Hint: work out the cost of the glass and frame separately.)

OpenStudy (anonymous):

20/h = A of base sqrt(20/h) = L 4*sqrt(20/h) = Perimeter of base h*4 = sum of sides 4*sqrt(20/h) = Perimeter of top 8*sqrt(20/h) + 4h = Sum of all twelve sides 10*(8*sqrt(20/h) + 4h) = 80sqrt(20/h) + 40h = Cost of entire metal frame 3*(4*h*sqrt(20/h) + (20/h))= Cost of glass 80sqrt(20/h) + 40h + 3*(4*sqrt(20h) + (20/h))= C

OpenStudy (anonymous):

got it? @butterflyprincess